The Chicago Bulls� arena is called the United Center and the Dallas Mavericks� arena is called the
American Airlines Center. They have a combined capacity of 42,700 people. Five times the
capacity of the American Airlines Center is 2,000 more than four times the capacity of the United
Center. What is the capacity of each arena?
Answer by mananth(16949) 
 The Chicago Bulls� arena is called the United Center and the Dallas Mavericks� arena is called the
American Airlines Center. They have a combined capacity of 42,700 people. Five times the
capacity of the American Airlines Center is 2,000 more than four times the capacity of the United
Center. What is the capacity of each are
They have a combined capacity of 42,700 people
Let the AA center hold x people
Five times the capacity of the American Airlines Center = 5x
2,000 more than four times the capacity of the United
Center. let UC hold y people
For equating
4y+2000
5x = 4y+2000
5x-4y=2000................1
Combined capacity = 42700
x+y=42700 ...................2
Multiply 2 by 4
4x+4y=170800 add ..........................3
add 2 & 3
9x = 172800
x =19,200
y is balance 23500
The capacity of the United Center is 23,500 and that of the American Airlines Center is 19,200.
check
5*19200-4*23500=2000
|
|
|